OpenStudy (atomicreaper):

What are the roots of -12x^2 + 11x − 3 = 0?

OpenStudy (unklerhaukus):

you could apply the quadratic formula

OpenStudy (atomicreaper):

explain

OpenStudy (unklerhaukus):

a quadratic equation \[ax^2+bx+c=0\] has roots \(x_1\), and \(x_2\) from the quadratic formula:\[x_{1,2} = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (unklerhaukus):

is this helpful?

OpenStudy (unklerhaukus):

if \(\sqrt{b^2-4ac}\) is less than zero, then there are no real roots